OpenStudy (anonymous):

you are making party invitations by hand. It has taken you 45 minutes to make 6 cards. If you continue at this rate, how many cards can you make in 2 hours?

jimthompson5910 (jim_thompson5910):

2 hours = 2*60 = 120 minutes

jimthompson5910 (jim_thompson5910):

(45 minutes)/(6 cards) = (120 minutes)/(x cards) 45/6 = 120/x 45*x = 6*120 I'll let you finish

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

I got 720

jimthompson5910 (jim_thompson5910):

6 times 120 is 720

jimthompson5910 (jim_thompson5910):

but that isn't the answer

jimthompson5910 (jim_thompson5910):

45 times _______ = 720 what goes in the blank? you could guess but to find out quickly, you would divide both sides by 45

jimthompson5910 (jim_thompson5910):

hopefully that makes sense

OpenStudy (anonymous):

oh, ok

jimthompson5910 (jim_thompson5910):

what answer do you get?